Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Enter a number. Round your answer to four decimal places.)
μ = 22; σ = 3.8
P(x ≥ 30) =

Answer by rothauserc(4718)

Probability (P) ( X > or = 30) = 1 - P( X < 30 )
:
z-score(30) = (30 - 22)/3.8 = 2.1053 is approximately 2.11
:
look in z-tables for the corresponding probability
:
P ( X > or = 30) = 1 - 0.9826 = 0.0174
: